# Fairness is Not Silence: Unmasking Vacuous Neutrality in Small Language Models

### v1

@@ Appendix A BBQ Dataset and Evaluation Metrics
To examine how model accuracy changes when constrained to provide unbiased answers in disambiguated examples, we compute a Bias Non-Alignment metric, which quantifies the impact of stereotype alignment on task performance. The evaluation set is partitioned into two subsets: Bias-Aligned, where the correct answer corresponds to the Target group, and Bias-Nonaligned, where it corresponds to the Non-Target group. For each model, the Bias Non-Alignment score is defined as the accuracy difference between bias-nonaligned and bias-aligned instances. Positive values indicate improved performance under bias rejection, suggesting that stereotype alignment previously hindered accuracy. Negative values suggest the opposite. This analysis helps distinguish genuinely fair models from those whose fairness may come at the cost of utility. Results are shown in Figure 6.
Appendix B Distribution of the predicted labels (A, B and C)
@@
In every BBQ instance, the three answer labels, A, B, and C, are dynamically shuffled, yet always map one-to-one onto the Target group (stereotype-consistent choice), the Non-Target group (counter-stereotypical choice), and the unknown option (indicating legitimate uncertainty). Because this mapping is randomized per question, the aggregate distribution of model selections across A, B, and C provides a sensitive diagnostic of positional bias: systematic overselection or avoidance of a given label suggests reliance on surface order rather than content. By comparing a model’s label frequencies with the ground truth proportions of the target, non-target, and unknown answers, we can unravel two complementary behaviors: vacuous neutrality and stereotypical alignment. A balanced distribution where selections of A, B, and C mirror their groundtruth prevalence across demographic categories signals robust handling of ambiguity and fair reasoning, whereas deviations from balance expose positional heuristics or unresolved biases that can undermine reliability in sensitive deployments.
Appendix C Compression Metrics
The metrics Relative F1 Change, Change in Bias, Compression Factor, and Shrinkage Ratio facilitate comparison of quantized models based on storage efficiency, bias reasoning performance, and deployment readiness.
Compression Factor:
Compression Factor measures how many times smaller a quantized model is compared to the original full-precision model:
Compression Factor=Original SizeCompressed $\text{Compression Factor}=\frac{\text{Original Size}}{\text{Compressed Size}}$ (5)
Shrinkage Ratio (%):
Shrinkage Ratio represents the relative reduction in model size due to compression:
[…]
Relative F1 Change:
Relative F1 Change reveals how much the model’s predictive accuracy changed after quantization, with positive values indicating improvement and negative values denoting performance loss. It is calculated as:
[…]
Change in Bias ($\Delta Bias$):
We compute the change in bias ($\Delta Bias$) score as a percentage-point change. Negative values indicate improved fairness (i.e., reduced bias), while positive values suggest an increase in bias post-compression.
$\Delta\text{Bias}=\text{Bias}_{\text{compressed}}-\text{Bias}_{\text{original}}$ (8)
@@ Appendix C Compression Metrics
Small Language Models Original Size Compressed Size Compression Factor Shrinkage Ratio (%)
LLaMA3.2-3B-Instruct 6.0 GB 2.2 GB 2.73$\times$ 63.3%
Qwen2.5-3B-Instruct 6.17 GB 2.69 GB 2.29$\times$ 56.4%
Phi-4-mini-Instruct 7.67 GB 2.8 GB 2.74$\times$ 63.5%
Table 3: Comparison of original and quantized SLM sizes using 4-bit AWQ compression.
@@
Figure 7: Impact of compression on model fairness and utility across bias categories. Each point reflects a model, context type, and bias category, positioned by the change in F1 score (x-axis) and change in bias score (y-axis) after compression.
Interpretation: Marker shapes distinguish models (circles: Phi-4-Mini, triangles: LLaMA3.2-3B, squares: Qwen2.5-3B); colors indicate context (blue: ambiguous, orange: disambiguated). The plot is divided into four quadrants: top-left ($x<0$, $y>0$) is harmful (worse accuracy and fairness), top-right is risky (better accuracy, worse fairness), bottom-left is acceptable (worse accuracy, better fairness), and bottom-right is ideal (better on both fronts).
LLaMA3.2-3B shows a favorable trend, many points lie below $y=0$, indicating reduced bias post-compression. Qwen2.5-3B’s points mostly lie near the origin, suggesting compression had little impact on performance or fairness. Phi-4-Mini shows mixed behavior, many points lie left of $x=0$, signaling performance drop, while the fairness impact is inconsistent, some categories improve (lower $y$), others worsen (higher $y$), highlighting Phi-4-Mini’s variable response to compression.

### v2

@@ A.1 Bias Non-Alignment
To examine how model competence changes when constrained to provide unbiased answers in disambiguated examples, we compute a Bias Non-Alignment metric, which quantifies the impact of stereotype alignment on task performance. The evaluation set is partitioned into two subsets: Bias-Aligned, where the correct answer corresponds to the Target group, and Bias-Nonaligned, where it corresponds to the Non-Target group. For each model, the Bias Non-Alignment score is defined as the accuracy difference between bias-nonaligned and bias-aligned instances. Positive values indicate improved performance under bias rejection, suggesting that stereotype alignment previously hindered accuracy. Negative values suggest the opposite. This analysis helps distinguish genuinely fair models from those whose fairness may come at the cost of utility. Results are shown in Figure 8.
A.2 Answer Choices {A, B, and C}
@@
In every BBQ instance, the three answer options {A, B, and C} are dynamically shuffled but maintain a one-to-one correspondence with the Target (stereotype-consistent), Non-Target (counter-stereotypical), and Unknown (legitimate uncertainty) labels. Because this mapping is randomized for each question, the aggregate distribution of a model’s selections across answer options serves as a sensitive diagnostic of positional bias: systematic preference or avoidance of a given label indicates reliance on positional heuristics rather than semantic reasoning. Comparing these label frequencies along with the ground-truth proportions of target, non-target, and unknown answers allows us to distinguish between two complementary behaviors - vacuous neutrality and stereotypical alignment. A balanced selection pattern, where model predictions approximate the true distribution across demographic categories and answer positions, reflects robust ambiguity handling and fair reasoning. Conversely, deviations from this balance reveal positional shortcuts or latent biases that undermine reliability in socially sensitive applications. The distribution of answer choices (A, B, C) across social categories can be seen in Figure 9 for Qwen2.5 family, Figure 10 for Llama3.2 family and Figure 11 for Gemma3 family. Table 5 summarizes the results for Small LMs (2B-4B), presenting their UR values, TNR values, distributions of choices over {A, B, C} and {S, AS, U}, and the corresponding Norm-DKL scores.
A.3 Evaluation Prompt & QA Instances
As shown in Figure 7, we display the evaluation prompt template used for SLMs (top) and representative BBQ examples from the Physical Appearance category (bottom) spanning different ambiguity and polarity settings. Each subfigure is a QA instance with three options {A, B, C} that correspond to Target, Non-Target, and Unknown; option positions are randomly shuffled and correct answers are boldfaced.
Tables 2, 3, and 4 present illustrative BBQ question pairs across all social bias categories. For each category, we include an ambiguous context (A) and its disambiguated counterpart (A+D), formed by combining implicit (A) and explicit (D) cues, along with a polarity pair, one negative (bias-reinforcing) and one non-negative (bias-negating). See the corresponding captions for interpretation details.
Appendix B StereoSet
StereoSet nadeem2020stereosetmeasuringstereotypicalbias is a bias evaluation dataset for language models that probes social stereotypes across categories such as Gender, Race Color, Religion and Socio Economic. In StereoSet, outputs are calculated based on the proportions of {S/AS/U} choices, where higher S than AS indicates stereotypical alignment, higher AS indicates counter-stereotypical preference, and U reflects abstention/irrelevance. The Stereo Score (SS) captures the tilt toward S vs. AS; the Language Modeling Score (LMS) measures preference for meaningful continuations (S or AS) over U; and the Idealized CAT Score (iCAT) combines SS and LMS to balance bias and utility.
SS (%) $=\frac{s}{s+as}\times 100,$ (5)
LMS $=\frac{s+as}{s+as+u}\times 100,$ (6)
iCAT $=\text{LMS}\times\frac{\min(\text{SS},\,100-\text{SS})}{50}.$ (7)
Appendix C CrowS-Pairs
CrowS-Pairs nangia2020crows is a minimal-pair bias benchmark in which each item contrasts a stereotypical and a anti-stereotypical sentence that differ only by a single, controlled lexical substitution, keeping topic and grammar fixed. Ground truth is specified at the level of polarity (stereo vs. anti-stereo) rather than a task-correct answer, which enables precise measurement of directional bias but does not, by design, assess utility or abstention. In our evaluation, we follow the StereoSet metrics, Stereo Score (SS), Language Modeling Score (LMS), and iCAT by mapping the stereotypical alternative to (S) and the anti-stereotypical alternative to (AS). To align calibration and ambiguity analysis with StereoSet, we extend CrowS-Pairs with a third “Unknown” (U) option, enabling unified reporting of SS, LMS, and iCAT and ensuring cross-benchmark comparability. We also shuffle option order and fix decoding settings to mitigate positional artifacts.
While StereoSet and CrowS-Pairs are informative for measuring directional social bias, they are not sufficient for assessing our framework: neither provides ground truth for task competence nor explicitly controls ambiguity (e.g., ambiguous vs. disambiguated contexts). Accordingly, we treat them primarily as reporting layers, reusing their Stereo Score (SS) and Language Modeling Score (LMS) and adding our $Norm-D_{KL}$ to probe positional bias, rather than as full evaluations of capability. Crucially, task competence remains unassessed: StereoScore is insensitive to the prevalence of the Unrelated ($U$) option, and LMS lacks external ground truth to verify correctness in QA-like settings. Thus, low bias scores on these datasets need not imply that a model is capable, calibrated, or useful under realistic ambiguity.
From Table  6 to Table  10, we report our zero-shot results on StereoSet and CrowS-Pairs for Small LMs (2B-4B). Because these datasets lack ground truth for task competence and do not provide explicit ambiguous or disambiguated contexts, we can only exercise Stage-1 of our framework, bias (e.g., the target/non-target ratio or StereoScore). While we can compute that ratio here, it merely replicates the Stage 1 signal and offers no evidence of task competence or calibrated ambiguity handling. Positional bias also cannot be meaningfully assessed, absent ground-truth positional labels, one can only compare to a uniform reference, which is uninformative. These limitations underscore the need for a complementary dataset that includes ambiguous situations with ground-truth answers for evaluating social biases more holistically, ideally, an additional BBQ-like resource with paired ambiguous/disambiguated contexts, per-item ground truth, and balanced label positions across social categories.
Appendix D Task Adaptation Finetuning
To examine how task adaptation influences reasoning and fairness, we fine-tuned all nine SLMs on CommonsenseQA (CSQA) talmor-etal-2019-commonsenseqa using parameter-efficient fine-tuning (PEFT) with LoRA adapters applied to attention and feedforward layers. We trained for 2 epochs using the AdamW optimizer with a cosine learning rate schedule and warmup, updating only adapter parameters while keeping the base model frozen. Training followed the multiple-choice QA format with a standard cross-entropy objective, and the same fixed train/validation data splits were used across all models for consistency. No fairness-oriented supervision or bias-mitigation losses were applied. After fine-tuning, models were directly evaluated on BBQ using the same multiple-choice prompting as in the main study to isolate how commonsense-oriented adaptation affects bias, task competence, positional bias and ambiguity handling. Table 1 presents evaluation results of all nine SLMs on the CommonsenseQA (CSQA) validation split. Overall, accuracy improves consistently with model scale across families, with Qwen2.5-3B and Phi-3.5-mini achieving the strongest performance. The results indicate that even SLMs demonstrate strong commonsense reasoning ability after task-specific fine-tuning while remaining computationally efficient.
As reported in the main text, this task-oriented adaptation substantially improves performance on disambiguated items while degrading reasoning under ambiguity across models, motivating Stages 3-4 of our framework.
Figure 5: Bias scores for CSQA-fine-tuned LMs on BBQ, shown as heatmaps for (a) Tiny LMs and (b) Small LMs under Ambiguous and Disambiguated contexts. Rows denote social bias categories and columns denote SLMs. Red indicates stereotypical, blue anti-stereotypical, and gray near-neutral responses. Most scores fall within -20% to +10%, with the range spanning –100% to +100%.
Table 1: Evaluation results of SLMs on the CommonsenseQA (CSQA) validation split.
Model Family Model Size Accuracy (Val)
Qwen 0.5B 0.676
1.5B 0.799
3B 0.838
LLaMA 1B 0.759
3B 0.823
Gemma 1B 0.694
4B 0.809
Phi 3.5B 0.834
4B 0.825
Figure 6: F1 scores for CSQA-fine-tuned LMs on BBQ, shown as heatmaps for (a) Tiny LMs (blue) and (b) Small LMs (green) under Ambiguous and Disambiguated contexts. Rows represent social bias categories and columns represent SLMs. Darker shades indicate higher F1 Score and stronger task performance; lighter shades denote weaker competence. Fine-tuned models show clear improvement, performing substantially better in disambiguated contexts but struggle in ambiguous contexts.
Bias Dimension Compared to the zero-shot setting in the main experiments, fine-tuning markedly increases bias in Tiny LMs up to about +20% while Small LMs remain near-balanced across categories. In Stage 1, bias magnitudes for Small LMs stay within ±10%, indicating that fine-tuning amplifies bias primarily in lower-capacity models, whereas larger ones retain stability and fairness (see Figure 5).
Utility Dimension We observe that both Tiny and Small Language Models perform strongly on disambiguated examples but fail substantially under ambiguous conditions. In particular, models such as LLaMA3.2-3B and Gemma3-4B achieve only single-digit F1 scores in ambiguous settings, while exceeding 90% on average across all social bias categories in disambiguated contexts. Even models that performed robustly in the main experiments, such as those from the Phi family, display the same pattern after fine-tuning. This sharp contrast indicates that, despite task-oriented adaptation, models remain brittle when reasoning under uncertainty, revealing persistent limitations in ambiguity handling despite strong overall competence in well-specified scenarios (see Figure 6).
Ambiguity Handling and Positional Bias Results for Stage 3 (Ambiguity Handling) and Stage 4 (Positional Bias) of Small LMs (2B-4B) are summarized in Table  11.